The st

9 The standard score for a two-tailed test is 2.0. What is the P-value? a. 0.9772 b. 0.0456 C. 0.0228 d. 97.72

I'm questioning my regarding the answer.   I'm stuck between 0.0456 and 0.0228. Can you please help.
Answer #1

Given that, the standard test score for a two-tailed test is 2.0. That is, test statistic is, z = 2.0

since, it is two-tailed test,

p-value = 2 * P(Z > 2.0) = 2 * 0.0228 = 0.0456

p-value = 0.0456